OpenStudy (anonymous):

find the solutions of the system y = x² + 3x – 4 y = 2x + 2

OpenStudy (anonymous):

(–3, 6) and (2, –4) (–3, –4) and (2, 6) (–3, –4) and (–2, –2) no solution

sam (.sam.):

y = x² + 3x – 4 y = 2x + 2 -------------------------------- x² + 3x – 4= 2x + 2 x^(2)+3x-4=2x+2 x^(2)+x-4=2 x^(2)+x-6=0 (x+3)(x-2)=0 x+3=0_x-2=0 x=-3,2

sam (.sam.):

(–3, –4) and (2, 6)

OpenStudy (ajprincess):

 Option B